# CoCoEvo: Co-Evolution of Programs and Test Cases to Enhance Code Generation

### v1

@@ III-C Test Case Evolution
Test fitness function. The evaluation of test cases involves two key metrics: confidence and discrimination. The confidence of test case is calculated as
where represents the total number of programs, represents the fitness of program , and is the evaluation matrix indicating whether passed . Intuitively, confidence measures the degree of agreement between the program population and the test cases, with higher program fitness contributing more weight to the degree of agreement.
@@
The selection of test cases is performed using the Pareto approach for multi-objective optimization [22], as illustrated in Fig. 5. The Pareto front is constructed based on the two optimization metrics, confidence and discrimination. Test cases with high confidence and low discrimination represent basic test cases that cover common scenarios, with a high probability of being correct. Test cases with high discrimination represent more challenging cases that are effective at exposing errors. We noticed that there are some test cases with very low confidence levels that sometimes have a very high degree of discrimination. To avoid this type of test case rom being selected, the test cases with confidence lower than the average in the selection results will be filtered out.
@@ III-D LLM-based evolution operators
Program crossover operator.
We utilize LLMs to perform program crossover, as illustrated in Fig. 6. The process involves the following steps:
@@
Information Construct. The prompt contains the following information: The entire existing population of test cases, the best program the program population and the line coverage information on it, where covered lines are marked as “[+]” and uncovered lines as “[-]”.
@@ IV-A Dataset
We collected a dataset of programming problems from the weekly and biweekly contests hosted on the LeetCode platform 111https://leetcode.com/, forming the LeetCode-Contest dataset. In the dataset, each problem includes a prompt comprising the function header and a natural language description in the form of a docstring, along with test cases and reference solutions. To ensure data integrity and avoid potential leakage, all selected problems were released after March 2024.
To more accurately assess the ability of LLMs to solve programming challenges, we took measures to prevent plagiarism of example test inputs and outputs in test case generation. Specifically, we removed all test input-output samples from the problem descriptions, retaining only the natural language problem statements, the function header, and the associated constraints.
For each problem, we collected Python solution codes and a substantial set of test cases available on the website. The detailed information of the LeetCode-Contest dataset is summarized in Table I.
| Problems | 80 |
| Average Test Cases | 644.40 |
| Selected Contests | weekly-contest-402, weekly-contest-401, biweekly-contest-132, weekly-contest-400, weekly-contest-399, biweekly-contest-131, weekly-contest-398, weekly-contest-397, biweekly-contest-130, weekly-contest-396, weekly-contest-395, biweekly-contest-129, weekly-contest-394, weekly-contest-393, biweekly-contest-128, weekly-contest-392, weekly-contest-391, biweekly-contest-127, weekly-contest-390, weekly-contest-389 |
@@ IV-B Models
We utilized four widely adopted and powerful models for our experiments: GPT-4o-mini [1], Qwen2.5-Coder-32B [3], Llama-3.1-70B [2], and DeepSeek-V3 [4]. For GPT-4o-mini, we accessed the API via the OpenAI platform 222https://platform.openai.com/. For Qwen2.5-Coder-32B and Llama-3.1-70B, we used the APIs available on the DeepInfra platform 333https://deepinfra.com/. For DeepSeek-V3, we utilized the API provided by the DeepSeek platform 444https://platform.deepseek.com/.
@@ IV-C Compared Baselines
Sampling [5]. A basic method that generates a large number of code samples and randomly selects one as the final answer.
@@
For all baselines, if the method requires test cases for evaluation or feedback generation, the test cases are generated using the same LLMs that perform code generation.
@@ IV-D Performance Indicators
The ground truth test cases in the dataset are used to evaluate the accuracy of the generated code. A code is considered correct if it passes all the test cases. The performance of the approaches on the dataset is measured using the pass@1 metric [46]. To calculate the pass@1 metric, each method submits a single program as the final answer for each problem. The pass@1 metric represents the percentage of problems successfully solved.
Additionally, for further analysis of the CoCoEvo approach, the ground truth program solutions are used to evaluate the correctness of the generated test cases. A generated test case is considered correct if it executes successfully on the ground truth solution.
@@ IV-E Experimental Settings
Given that invoking LLMs for code generation or repair consumes significantly more time and resources compared to executing code, we evaluated the performance of all methods under the condition that the number of LLM calls for code generation or repair is held constant. To ensure practical usability and align with the parameter settings of most code generation methods, we set the number of LLM calls for code generation to iterations. For methods requiring LLM-generated test cases, we limited test case generation to iterations, with a maximum of test cases extracted per iteration.
The parameter settings of the CoCoEvo method can be found in Table II.
| Program Population Size | 10 |
| Iterations | 10 |
| Min Test Case Population Size | 10 |
We observed that when calculating the pass@1 metric, if multiple codes achieve the same score, one code will be randomly selected as the final result. This introduces a certain degree of randomness to the evaluation process. To mitigate this, for all methods, we independently and randomly repeated the “selection evaluation” process times and reported the averaged results.
@@ V-A Comparison of Accuracy
In this section, we present the results of the accuracy comparison for the methods. Table III reports the pass@1 metric for each method across different LLMs on the LeetCode-Contest dataset. As shown in Table III, our proposed CoCoEvo method achieves the highest performance across all LLMs, demonstrating its superior ability to generate accurate code. This success can be attributed to the co-evolution framework employed by CoCoEvo, which facilitates a more effective iterative refinement between programs and test cases.
An interesting observation concerns the repair-based methods such as Self-Repair, Reflexion, and INTERVENOR. When these methods utilize test cases generated by LLMs instead of pre-defined ones, their effectiveness diminishes significantly. In some instances, their performance is even worse than that of the basic Sampling approach. This is because the erroneous test cases generated by LLMs can mislead these methods by providing inaccurate feedback, causing the repair process to make suboptimal adjustments to the code. Consequently, this results in lower accuracy and poorer overall performance. Similarly, for methods like Sampling+Filtering, where test cases are used to filter and select the best code, performance remains comparable to the Sampling method when the test cases are generated by LLMs. This highlights a critical limitation: these methods struggle to adapt effectively in scenarios where reliable pre-defined test cases are unavailable.
Furthermore, methods that leverage the agreement between programs and test cases, such as MBR-Exec and CodeT, exhibit higher performance compared to Sampling and repair-based approaches. Among these, CodeT stands out, delivering consistent improvements in pass@1 across all LLMs. This suggests that considering the alignment and mutual validation of programs and test cases is a promising strategy for enhancing code generation accuracy.
@@
Fig. 10 illustrates the changes in the accuracy of the test population throughout the evolution process. At initialization, the CoCoEvo method achieves a slightly higher test accuracy than the randomly generated method. Specifically, the CoCoEvo method generates test cases once and extracts test cases, whereas the random method involves generating test cases times and extracting test cases each time. This indicates that as the number of test case generations by LLMs increases, test accuracy tends to decrease slightly.
Interestingly, For the CoCoEvo method, the peak test accuracy is observed in the second iteration. In subsequent iterations, the accuracy declines slightly, potentially due to the Pareto selection process, where test cases with higher discrimination but lower confidence are accepted. This may inadvertently introduce some incorrect test cases. Nevertheless, the decline is gradual, and the test accuracy remains consistently high, making this behavior acceptable.
@@ V-B Comparison of Program Evolution Scheduler
| Crossover Rate Scheduler | Pass@1 |
| Constant | 51.00 |
| Cosine | 55.75 |
To assess the effectiveness of the crossover rate scheduler, we compare the cosine scheduler with a constant scheduler in our experiments, using a mutation rate of and a crossover rate of . These experiments were performed on the Qwen2.5-Coder-32B and the LeetCode-Contest dataset. The results, summarized in Table IV and Figure 11, highlight key differences between the cosine scheduler and the constant rate approach.
At the early stages of the evolution process, the constant method, characterized by a fixed crossover and mutation rate, demonstrates faster convergence and initially outperforms the cosine scheduler due to its higher crossover rate. However, this rapid convergence causes the Constant method to stagnate and become trapped in a local optimum, ultimately limiting its overall performance.
In contrast, the cosine scheduler, despite its lower initial performance, maintains consistent improvement throughout the evolution process. By adapting the crossover rate dynamically, the scheduler avoids premature convergence and achieves superior results in later stages. Notably, it surpasses the performance of the Constant method after the fifth iteration, ultimately delivering a higher pass@1 rate.
This comparison underscores the importance of dynamic scheduling in optimizing program evolution, particularly in scenarios prone to local optima.
@@ V-C Comparison of Program Fitness Function
| Program Fitness Function | Pass@1 |
| Pass Rate | 53.50 |
| CodeT Score | 55.75 |
To evaluate the impact of different program fitness functions, we conducted a comparative analysis using the CodeT score and the simple pass rate as fitness functions. The base model for these experiments was Qwen2.5-Coder-32B with the LeetCode-Contest dataset serving as the benchmark. The results are summarized in Table V and illustrated in Figure 12.
The findings reveal the insight that calculating the pass rate on test cases directly as the fitness function yields a lower pass@1 value compared to utilizing the CodeT score. This behavior can be attributed to the advantages of the CodeT score, which incorporates the congruence between programs and test cases more effectively. By accounting for this alignment, the CodeT score mitigates the negative effects of erroneous or poorly designed test cases on program fitness, leading to more stable and improved performance. These results underscore the superiority of the CodeT score as a program fitness function, especially in scenarios where robustness and convergence are critical.
@@ V-D Comparison of Test Fitness Function
| Test Fitness Function | Pass@1 |
| Failure Rate | 37.00 |
| Pass Rate | 51.50 |
| Confidence | 51.50 |
| Pareto | 55.75 |
In this section, we present the results of comparing various test fitness functions. Specifically, we evaluate Failure Rate, Pass Rate, Confidence, and Pareto as methods for test fitness calculation. The experiments are conducted using the Qwen2.5-Coder-32B model on the LeetCode-Contest dataset. The detailed results are depicted in Fig. 13, and additional information on test case accuracy is provided in Fig. 14.
@@
Our findings indicate that employing the Pareto method as the selection mechanism yields the best performance. The Pareto approach prioritizes test cases with both high discrimination and high confidence levels, thereby demonstrating superior effectiveness. In contrast, using Failure Rate as the fitness function results in a sharp decline in pass@1 performance. Previous studies on co-evolution have generally argued that the confidence level of a test case increases as more programs fail on it. However, we discovered that if the test cases contain errors, this method can lead to a population dominated by faulty test cases, which fail to provide accurate feedback to the programs.
@@
Furthermore, Fig. 13 reveals that employing Confidence as the test fitness function outperforms using Pass Rate by a small margin. The Confidence function weights the fitness of test cases based on program performance, with programs demonstrating higher fitness values contributing more heavily to the confidence score. This approach ensures that more reliable programs have a greater influence on test case evaluation.
V-E Ablation Study

### v2

@@ III-C Test Case Evolution
Test case evaluation. The evaluation of test cases involves two key metrics: Confidence and Discrimination. Confidence evaluates the probability that a test case is correct, and discriminative power evaluates a test case’s ability to distinguish between programs. The Confidence of test case is calculated as
where represents the size of the program population, represents the fitness of program , and is the evaluation matrix indicating whether passed . Intuitively, Confidence measures the degree of agreement between the program population and the test cases, with higher program fitness contributing more weight to the degree of agreement.
@@
Test case selection. The selection of test cases is performed using the Pareto approach for multi-objective optimization [22]. The Pareto frontier, also known as the Pareto front or Pareto boundary, represents the set of non-dominated solutions in a multi-objective optimization problem. A solution is considered Pareto optimal if no other feasible solution improves one objective without causing a deterioration in at least one other. Formally, given a vector-valued objective function , a solution is Pareto optimal if there does not exist another solution such that for all , with at least one strict inequality.
In test case selection, the Pareto front is constructed based on the two metrics, and , which represent the Confidence of test cases and the Discrimination of test cases. The Pareto front is illustrated in Fig. 5. Test cases with high and low represent basic test cases that cover common scenarios, with a high probability of being correct. Test cases with high represent more challenging cases that are effective at exposing errors. During the selection process, preference is given to individuals that perform well on both indicators. This strategy aims to: (1) preserve the correctness of the test case population, ensuring accurate evaluation of the program, and (2) retain test cases with high discriminative power to maintain selection pressure on the program population. We also noticed that there are some test cases with very low that sometimes have a relatively high . To avoid adverse effects of these test cases on program evolution, the test cases with lower than the average in the selection results will be filtered out.
@@ III-D LLM-based Evolution Operators
Program crossover operator. We utilize LLMs to perform program crossover, as illustrated in Fig. 6. The process involves the following steps:
@@
Information Construct. The prompt contains the following information: The entire existing population of test cases, the best program in the program population, and the line coverage information on it, where covered lines are marked as “[+]” and uncovered lines as “[-]”.
@@ IV-A Dataset
We collected a dataset of programming problems from the weekly and biweekly contests hosted on the LeetCode platform 111https://leetcode.com/, forming the LeetCode-Contest dataset. In the dataset, each problem includes a prompt comprising the function header and a natural language description in the form of a docstring, along with program solutions and approximately ground truth test cases per problem. These comprehensive ground truth test cases can closely simulate the real-world scenario of submissions on LeetCode, making our evaluation more convincing. To ensure data integrity and avoid potential leakage, all selected problems were released after March 2024. To more accurately assess the ability of LLMs to solve programming challenges, we took measures to prevent plagiarism of example test inputs and outputs in test case generation. Specifically, we removed all test input-output samples from the problem descriptions, retaining only the natural language problem statements, the function header, and the associated constraints. The detailed information of the LeetCode-Contest dataset is summarized in Table I.
Problems | 80 | ||||||||||
Average Test Cases | 644.40 | ||||||||||
Selected Contests |
@@ IV-B Models
We utilized four widely adopted and powerful LLMs for our experiments: GPT-4o-mini [1], Qwen2.5-Coder-32B [3], Llama-3.1-70B [2], and DeepSeek-V3 [4]. For GPT-4o-mini, we accessed the API via the OpenAI platform 222https://platform.openai.com/. For Qwen2.5-Coder-32B and Llama-3.1-70B, we used the APIs available on the DeepInfra platform 333https://deepinfra.com/. For DeepSeek-V3, we utilized the API provided by the DeepSeek platform 444https://platform.deepseek.com/.
@@ IV-C Compared Baselines
Sampling. A basic method that generates a large number of code samples and randomly selects one as the final answer.
@@
For the baselines (Sampling+Filtering, Self-Repair, Reflexion, INTERVENOR) that originally relied on pre-defined test cases for program evaluation, we adapt them to our setting by replacing the pre-defined test cases with LLM self-generated test cases.
@@ IV-D Performance Indicators
The ground truth test cases in the dataset are used to evaluate the accuracy of the generated program. A code is considered correct if it passes all the ground truth test cases. The performance of the approaches on the dataset is measured using the pass@1 metric [46]. To calculate the pass@1 metric, each method submits a single program as the final answer for each problem. The pass@1 metric represents the percentage of problems successfully solved. Additionally, for further analysis of the CoCoEvo approach, the ground truth program solutions are used to evaluate the correctness of the generated test cases. A generated test case is considered correct if it executes successfully on the ground truth solution.
@@ IV-E Experimental Settings
Following the comparison settings used in prior studies [7, 13, 12], we compare the performance of different methods by constraining the number of code generation attempts made by the LLMs to be equal. Based on the common limitations of these approaches and taking into account the cost of invoking LLMs, we limit the number of LLM code generation calls to . For methods that require test case generation, we impose a similar constraint by limiting the number of test case generation calls made to the LLMs to , with a maximum of test cases extracted per generation. For CoCoEvo, we adopt a balanced configuration by setting both the population size and the number of iterations to . Comparative experiments with alternative population sizes and detailed parameter settings of baselines and CoCoEvo can be found in the supplementary material. Since different code generation strategies may require varying numbers of tokens, we provide a detailed analysis of token consumption in Section V-G. We observed that when calculating the pass@1 metric, if multiple codes achieve the same score, one code will be randomly selected as the final result. This introduces a certain degree of randomness to the evaluation process. To mitigate this, for all methods, we independently and randomly repeated the “selection evaluation” process times and reported the averaged results.
@@ V-A Comparison of Accuracy
In this section, we present the results of the accuracy comparison for the methods. Table II reports the pass@1 metric for each method across different LLMs on the LeetCode-Contest dataset. As shown in Table II, our proposed CoCoEvo method achieves the highest performance across all LLMs, demonstrating its superior ability to generate accurate code. This success can be attributed to the co-evolution framework employed by CoCoEvo, which facilitates a more effective iterative refinement between programs and test cases.
An interesting observation concerns the code repair based methods such as Self-Repair, Reflexion, and INTERVENOR. When these methods utilize test cases generated by LLMs instead of pre-defined ones, their effectiveness diminishes significantly. In some instances, their performance is even worse than that of the basic Sampling approach. This is primarily due to the faulty LLM-generated tests mislead the repair process, leading to suboptimal code modifications and degraded overall performance. Similarly, Sampling+Filtering yields only a slight improvement over plain Sampling under these conditions, highlighting these approaches’ dependence on reliable test suites.
Furthermore, methods that leverage the agreement between programs and test cases, such as MBR-Exec and CodeT, outperform both Sampling and repair-based methods. Among these, CodeT stands out, delivering consistent improvements in pass@1 across all LLMs, suggesting that mutual validation of code and test cases is a promising strategy for improving generation accuracy.
@@
Fig. 10 presents the accuracy trends of the test case population and offspring over the evolutionary iterations, as well as the accuracy of the randomly generated test cases. As shown in the figure, the accuracy of the test case population consistently exceeds that of randomly generated test cases, demonstrating the effectiveness of the alternating selection strategy in the co-evolutionary framework. This indicates that selecting test cases based on their Confidence is beneficial for identifying faulty test cases. Moreover, the strategy for generating additional test cases exhibits a modest upward trend in the accuracy of offspring test cases, which also contributes positively to maintaining the overall accuracy of the test case population.
Interestingly, the peak test accuracy of the test population is observed in the second iteration. In subsequent iterations, the accuracy declines slightly. The slight decline can be attributed to the presence of a few incorrect test cases on the Pareto front that exhibit high discriminative capability but low Confidence, thus affecting overall accuracy. Nevertheless, the Pareto selection approach preserves a substantial number of accurate and highly discriminative test cases, thereby offering greater benefits to the co-evolutionary process. Consequently, the marginal reduction in accuracy remains within an acceptable range.
@@ V-B Comparison of Program Evolution Scheduler
Model | Crossover Rate | Pass@1 |
GPT-4o-mini | Constant | 45.00 |
Cosine | 49.75 | |
Qwen2.5-Coder-32B | Constant | 51.00 |
Cosine | 55.75 | |
Llama-3.1-70B | Constant | 42.75 |
Cosine | 45.00 | |
DeepSeek-V3 | Constant | 72.50 |
Cosine | 76.25 |
To assess the effectiveness of the crossover rate scheduler, we compare the cosine scheduler with a constant scheduler in our experiments, using a mutation rate of and a crossover rate of . These experiments were performed on the four LLMs and the LeetCode-Contest dataset. The results, summarized in Table III and Figure 11, highlight key differences between the cosine scheduler and the constant rate approach.
At the early stages of the evolution process, the constant method, characterized by a fixed crossover and mutation rate, demonstrates faster convergence and initially outperforms the cosine scheduler on Qwen2.5-Coder-32B, Llama-3.1-70B, and DeepSeek-V3. However, this rapid convergence causes the Constant method to stagnate and become trapped in a local optimum, ultimately limiting its overall performance. In contrast, the cosine scheduler, despite its lower initial performance, maintains consistent improvement throughout the evolution process. On GPT-4o-mini, the cosine scheduler consistently outperforms the method with a constant crossover rate.
By adapting the crossover rate dynamically, the scheduler avoids premature convergence and achieves superior results in later stages. Notably, it surpasses the performance of the Constant method after the fifth iteration, ultimately delivering a higher pass@1 rate. This comparison underscores the importance of dynamic scheduling in optimizing program evolution, particularly in scenarios prone to local optima.
@@ V-C Comparison of Program Fitness Function
Model | Program Fitness Function | Pass@1 |
GPT-4o-mini | Pass Rate | 45.50 |
CodeT Score | 49.75 | |
Qwen2.5-Coder-32B | Pass Rate | 53.50 |
CodeT Score | 55.75 | |
Llama-3.1-70B | Pass Rate | 41.00 |
CodeT Score | 45.00 | |
DeepSeek-V3 | Pass Rate | 73.25 |
CodeT Score | 76.25 |
To evaluate the impact of different program fitness functions, we conducted a comparative analysis using the CodeT score and the simple pass rate as fitness functions. These experiments are performed on the four LLMs and the LeetCode-Contest dataset. The results are summarized in Table IV and illustrated in Figure 12.
The results indicate that using the raw pass rate as the fitness function yields a lower pass@1 than the CodeT score. This is because CodeT better captures the alignment between programs and test cases, reducing the impact of flawed or poorly designed tests on fitness. Consequently, CodeT delivers more stable performance and faster convergence. These findings highlight CodeT’s superiority as a fitness function, particularly when robustness and convergence are essential.
@@ V-D Comparison of Test Fitness Function
Model | Test Fitness Function | Pass@1 |
GPT-4o-mini | Failure Rate | 37.50 |
Pass Rate | 47.00 | |
Confidence | 45.50 | |
Pareto | 49.75 | |
Qwen2.5-Coder-32B | Failure Rate | 37.00 |
Pass Rate | 51.50 | |
Confidence | 51.50 | |
Pareto | 55.75 | |
Llama-3.1-70B | Failure Rate | 26.00 |
Pass Rate | 37.50 | |
Confidence | 36.75 | |
Pareto | 45.00 | |
DeepSeek-V3 | Failure Rate | 66.75 |
Pass Rate | 71.75 | |
Confidence | 72.25 | |
Pareto | 76.25 |
In this section, we present the results of comparing various test fitness functions. Specifically, we evaluate Failure Rate, Pass Rate, Confidence, and Pareto as methods for test fitness calculation. These experiments are performed on the four LLMs and the LeetCode-Contest dataset. The detailed results are depicted in Fig. 13, and additional information on test case accuracy is provided in Fig. 14.
@@
Experimental results indicate that employing the Pareto method as the selection mechanism yields the best performance. In contrast, using Failure Rate as the fitness function results in a sharp decline in pass@1 performance. Previous studies on co-evolution have generally argued that the strength of a test case increases as more programs fail on it. However, we discovered that if the test cases contain errors, this method can lead to a population dominated by faulty test cases, which fail to provide accurate feedback to the programs.
@@
Furthermore, Fig. 13 reveals that employing Confidence as the test fitness function slightly outperforms using Pass Rate in most cases. By weighting test cases according to program performance, higher-performing programs contribute more to the Confidence score, ensuring that the most reliable programs have the greatest influence on evaluation.
V-E Analysis of Test Coverage Impact
Sampling+Filtering | 37.50 | 44.50 | 31.00 | 68.25 |
Sampling+Filtering w/ cov | 36.25 | 47.75 | 33.25 | 68.50 |
MBR-Exec | 33.75 | 45.00 | 38.75 | 70.00 |
MBR-Exec w/ cov | 35.00 | 45.00 | 33.75 | 72.50 |
CodeT | 46.25 | 47.50 | 41.25 | 72.50 |
CodeT w/ cov | 47.75 | 50.25 | 39.00 | 71.50 |
CoCoEvo w/o cov | 48.25 | 53.00 | 42.75 | 76.00 |
CoCoEvo | 49.75 | 55.75 | 45.00 | 76.25 |
During offspring test case generation, we incorporate coverage information of the current test case population on the best-performing program individual to guide the generation of more targeted test cases. To explore whether this strategy could improve baseline performance, we applied it to several compatible baselines. Since this strategy requires an initial set of programs and then iteratively generates test cases based on their coverage, we adopt this method for Sampling, Sampling+Filtering [5], MBR-Exec [13], and CodeT [12]. We also conducted experiments to evaluate the performance of CoCoEvo without the coverage information. The results are shown in Table VI.
The experiments reveal that simply adding test coverage information to baseline methods can sometimes negatively affect performance, as observed in GPT-4o-mini with Sampling+Filtering, Llama-3.1-70B with MBR-Exec and CodeT, and DeepSeek-V3 with CodeT. In contrast, CoCoEvo benefited consistently from coverage information across all LLMs, underscoring that such data is most effective when used in conjunction with the co-evolutionary framework.
V-F Ablation Study
Model | Method | Pass@1 |
GPT-4o-mini | Sampling | 35.25 |
w/o test evolution | 48.00 | |
CoCoEvo | 49.75 | |
Qwen2.5-Coder-32B | Sampling | 44.00 |
w/o test evolution | 51.75 | |
CoCoEvo | 55.75 | |
Llama-3.1-70B | Sampling | 32.00 |
w/o test evolution | 43.00 | |
CoCoEvo | 45.00 | |
DeepSeek-V3 | Sampling | 69.00 |
w/o test evolution | 74.50 | |
CoCoEvo | 76.25 |
We conduct ablation experiments across the four LLMs on the LeetCode-Contest dataset to evaluate the effectiveness of the test case evolution module. In these experiments, we removed the test case evolution module and directly utilized the test cases generated by the LLMs to evaluate the code. The comparison results are presented in Table VII and Fig. 15.
The experimental results demonstrate that using only the program evolution module, without incorporating the test case evolution module, still outperforms the Sampling baseline. However, after integrating the test case evolution module, the performance of CoCoEvo is further enhanced. This highlights the critical role of the test case evolution module in the co-evolution process.
V-G Analysis of Token Usage
Method | GPT-4o-mini | Qwen2.5-Coder-32B | Llama-3.1-70B | DeepSeek-V3 | |||||||
Prompt | Completion | Prompt | Completion | Prompt | Completion | Prompt | Completion | ||||
Sampling | 25823.75 | 14326.96 | 26088.75 | 15082.05 | 25900.00 | 13754.51 | 25821.25 | 19535.79 | |||
Sampling+Filtering | 28767.25 | 18115.60 | 28932.63 | 18904.11 | 28859.50 | 16305.21 | 28763.38 | 25686.94 | |||
Self-Repair | 40383.83 | 23086.20 | 40033.05 | 25913.49 | 39661.24 | 23284.41 | 41461.88 | 32701.89 | |||
Reflexion | 221277.83 | 48270.31 | 236637.13 | 83716.98 | 234549.55 | 54254.50 | 238286.46 | 57093.93 | |||
INTERVENOR | 107996.59 | 79492.75 | 105999.59 | 93253.63 | 131700.33 | 84377.99 | 102779.86 | 85323.98 | |||
CodeCoT | 44524.69 | 20815.48 | 49635.12 | 19936.05 | 38194.79 | 22017.59 | 41806.95 | 18614.76 | |||
AgentCoder | 73915.88 | 35099.35 | 81683.24 | 33976.43 | 71640.85 | 39185.11 | 62136.31 | 27798.86 | |||
MBR-Exec | 28767.25 | 18115.60 | 28932.63 | 18904.11 | 28859.50 | 16305.21 | 28763.38 | 25686.94 | |||
CodeT | 28767.25 | 18115.60 | 28932.63 | 18904.11 | 28859.50 | 16305.21 | 28763.38 | 25686.94 | |||
CoCoEvo Stop-1 | 13558.04 | 5706.66 | 11414.38 | 5404.08 | 15792.08 | 7535.93 | 10031.14 | 5538.18 | |||
CoCoEvo Stop-2 | 20403.58 | 8255.95 | 17998.85 | 8118.11 | 23626.14 | 10699.68 | 15988.85 | 8192.44 | |||
CoCoEvo Stop-3 | 26739.18 | 10559.31 | 25055.68 | 11095.44 | 30715.03 | 13666.15 | 22749.20 | 11127.73 | |||
CoCoEvo Stop-4 | 33405.58 | 12959.24 | 32192.81 | 14077.55 | 38130.99 | 16786.09 | 30061.74 | 14264.30 | |||
CoCoEvo Stop-5 | 40120.51 | 15358.71 | 39038.18 | 16813.46 | 44900.06 | 19619.78 | 37406.18 | 17349.59 | |||
CoCoEvo | 61567.61 | 22947.25 | 63638.83 | 26393.06 | 60122.26 | 26128.68 | 66190.20 | 28926.00 |
CoCoEvo Stop-1 | 47.75 | 48.75 | 39.75 | 74.75 |
CoCoEvo Stop-2 | 47.50 | 51.50 | 39.50 | 76.00 |
CoCoEvo Stop-3 | 46.00 | 51.00 | 42.75 | 75.25 |
CoCoEvo Stop-4 | 48.00 | 54.00 | 44.00 | 76.00 |
CoCoEvo Stop-5 | 48.75 | 54.00 | 43.00 | 75.25 |
CoCoEvo | 49.75 | 55.75 | 45.00 | 76.25 |
Some strategies, such as reflection and repair employed in some baseline methods, as well as the crossover, mutation, and additional test case generation strategies used in CoCoEvo, may consume more tokens than others. We therefore report each method’s token consumption to compare their resource efficiency. Table VIII shows the average token consumption per problem on the LeetCode-Contest dataset. We report input (Prompt) and output (Completion) tokens separately, as output tokens are generally more costly. Notably, we designed an early stopping strategy to investigate whether the token consumption of the CoCoEvo method could be optimized. Since problem difficulty varies, simpler problems often require fewer iterations, whereas more complex problems tend to require more. Thus, the early stopping mechanism is triggered when the code population maintains the same fitness value and passes the same test cases for consecutive generations, indicating convergence. In Table VIII, we report token consumption for different values of (ranging from to ), denoted as CoCoEvo Stop-1 to CoCoEvo Stop-5, respectively.
As shown in Table VIII, Sampling consumes the fewest tokens because it only generates code, not test cases. Meanwhile, Sampling+Filtering [5], MBR-Exec [13], and CodeT [12] show identical token consumption as they share the same generated programs and test cases, differing only in their final program selection logic. Code repair methods (e.g., Self-Repair [7], Reflexion [6], INTERVENOR [8], CodeCoT [12], and AgentCoder [16]) consume significantly more tokens because their prompts must include the original code and test execution results. Reflexion consumes the most input tokens because it first generates a reflective analysis before rewriting the code. CoCoEvo’s input token usage is higher than CodeT’s because its crossover, mutation, and test generation modules require existing programs and test cases as input prompts. However, its output token usage is lower than that of repair-based methods.
Notably, we observe that with the integration of the early stopping strategy, the token consumption of CoCoEvo is significantly reduced. The required input tokens become comparable to those of CodeT, while the output token usage is lower. To examine the impact of the early stopping strategy on accuracy, we report the pass@1 performance of CoCoEvo on the LeetCode-Contest dataset under different values of . The results are presented in Table IX. We find that, when jointly considering token consumption and accuracy, setting the early stopping parameter to leads to a slight decrease in accuracy compared to configurations without early stopping, but achieves a 30% to 50% reduction in overall token usage, demonstrating the considerable potential of CoCoEvo.
V-H Evaluation on Real-World Projects
Project | Link | Problems |
mrjob | https://github.com/Yelp/mrjob | 1 |
Jinja2 | https://github.com/pallets/jinja | 1 |
twilio-fatisar | https://github.com/twilio/twilio-python | 1 |
feedparser | https://github.com/kurtmckee/feedparser | 1 |
dash | https://github.com/plotly/dash | 2 |
diffprivlib | https://github.com/IBM/differential-privacy-library | 3 |
databases | https://github.com/encode/databases | 1 |
This section evaluates CoCoEvo’s performance on real-world software projects. We selected problems from the DevEval dataset [47], where each problem consists of a function and its corresponding unit tests extracted from a real-world project. To ensure realistic and manageable problems, we filtered for functions with: (1) at least one external dependency, (2) five or more unit tests, and (3) to lines of code. In total, we selected problems from software projects, as summarized in Table X.